The pressure of a sample of helium in a 1.0L container is 0.98 atm. What is the new pressure if the sample is placed in a 2.0
L container?
Law being used: ___________ FORMULA WORK ANSWER
Given 1:__________
Given 2:__________
Given 3:__________
1
Expert's answer
2020-04-18T07:00:30-0400

Solution:

Temperature is constant


Law being used: Boyle’s Law


Given 1: __ P1 = 0.98 atm __

Given 2: __ V1 = 1.0 L __

Given 3: __V2 = 2.0 L __


FORMULA:

P1V1 = P2V2


WORK:

P1 * V1 = P2 * V2

(0.98 atm) * (1.0 L) = (P2) * (2.0 L)

P2 = (0.98 * 1.0) / 2.0 = 0.49 atm

P2 = 0.49 atm


ANSWER: P2 = 0.49 atm.